assume that an investor is willing to pay $908.32 for a bond (pv 1000, coupon rate 8%, maturing in 20 years)

what is the investor's required rate of return, kd?

(bond issuer's viewpoint) what if the net price after flotation costs is $850, what then will kbe?

what is the after-tax kassuming tax rate of 30%?



1
Expert's answer
2021-10-03T14:17:20-0400

a)

 present value of annuity:

"P\frac{1-(1+k_d)^{-n}}{k_d}"

where P is periodic payment,

kd is rate of return,

n is number of periods.


 present value of the bond:

"P\frac{1-(1+k_d)^{-n}}{k_d}+1000\cdot (1+k_d)^{-n}"


"80\cdot\frac{1-(1+k_d)^{-20}}{k_d}+1000\cdot (1+k_d)^{-20}=908.32"

Using Bisection mehtod on online calculator https://atozmath.com/, we get:

"k_d=0.09=9\%"

b)

"80\cdot\frac{1-(1+k_d)^{-20}}{k_d}+1000\cdot (1+k_d)^{-20}=850"

"k_d=0.0973=9.73\%"


c)

after-tax rate:

"0.08\cdot (1-0.3)=0.056=5.6\%"